• Welcome to ADA Depot - A Forum To Support Users of ADA Amplification Gear.
 

News:

Lets get Technical > All Things Tube All about preamp and power amp Tubes.

Main Menu

MP-2 S/N 501293 - haunted?

Started by Gregg, July 01, 2024, 04:01:23 PM

Previous topic - Next topic

Gregg

Note: I already reached out to MarshallJMP privately.  He responded super quickly suggesting doing a public post to get more thoughts on this one.

Hello folks -

I have been working on an MP-2 that seems borderline haunted at this point.  Below are my repair attempt notes.  I have included links to some marked up schematics that highlight parts I removed, swapped, or tested.  Also clarified some of the blurry text.

Overall, I have made some progress in that things are less intermittent, but am starting to doubt my own sanity.  General observation about this MP-2 501293 is that there were more poor solder joints and signs of rework (flux) on the PCB than other MP-2s PCBs I have seen.

Any suggestions would be appreciated.  I'm determined to exorcise whatever ghosts / demons haunt this unit.

Gregg



Marked up schematics & pictures

Problem
Unit had unspecified issues per owner.  Never really worked right when he purchased it back in the day on eBay.  When warm it would sometimes work correctly.

Repair Log
Performed MP2 noise mod and noise gate mod per ADA Forum with DigiKey sourced parts.  All voltages including HV are within spec and without ripple.

No output when cold.  Once warm it would work OK for a while, but then fail.
  • Replaced Q200 & Q202 (2SK364)
  • Replaced Q203 & Q204 (2SK364)

Thermal issue resolved.

  • Replaced Q306 (2N2907) level shifter for WAH +
  • Tested D209 & D210 - appear OK
  • Check voltages at D206 & D208 - appears OK
  • Reflowed some bad solder joints.

Output from MP2B @R242 goes away when either WAH is IN or Compressor is OUT.  Testing with a 1 khz .5 vpp signal from a signal generator.

  • Swapped the SSM2120 (U210) - no change
  • Swapped OPA2604AP U207 and U205 - no change
  • Presence of U205 and U207 - no difference (???)
  • Removed C205 and C206 to eliminate coupling from the WAH  - no change
  • WAH IN should kill output
  • Lifted R595 to disconnect the DOOR_SLAMMER circuit - no change
  • Removed C230 - couples the input signal to the switching transistor Q200.  C230 tests OK out of circuit. Input signal @ R295 good even when toggling WAH or Compressor IN/OUT.
  • Reinstalled C230.
  • Planned
    • Reinstall C205 and C206
    • Observe signal @ R240 (100K).  Thinking there is an issue with Q200 (2SK364), despite having replaced it, or the level shifter.  Level shifter parts: D206 (1N4148), D208 (1N4148), Q220 (2N2907), Q221 (2N3904)